Required information [The following information applies to the questions displayed below.) Franklin Company is a retail compaSalary expense (fixed) Sales commissions Supplies expense Utilities (fixed) Depreciation on store fixtures (fixed)* Rent (fixComplete this question by entering your answers in the tabs below. Required A Required B Required C Required D Required E ReqThe cost of goods sold is 60 percent of sales. The company desires to maintain a minimum ending inventory equal to 10 percentPrepare a selling and administrative expenses budget. November December October Selling and Administrative Expense Budget SalUtilities and sales commissions are paid the month after they are incurred; all other expenses are paid in the month in whichFranklin borrows funds, in increments of $1,000, and repays them on the last day of the month. Repayments may be made in any
Answer #1


Franklin Company Sales Budgets Cash Sales Credit Sales Total Budgeted Sales October November December 72000 86400 103680 1080

Selling and Administrative Expense Budget October November December Total Salary Expense 18600 18600 18600 55800 Sales Commis

A 1 Franklin Company 3 Sales Budgets 5 Cash Sales 6 Credit Sales 7 Total Budgeted Sales October =B7*0.4 =B7-B5 180000 Novembe